OpenStudy (anonymous):

What is the vertex of the function f(x) = 7x2 - 28x - 60? (2, -88) (3, 56) (2, 88) (3, -56) confuse...

Mehek (mehek14):

use \(\dfrac{-b}{2a}\)

OpenStudy (anonymous):

-28 b?

OpenStudy (anonymous):

and 7 a?

Mehek (mehek14):

\(\dfrac{-(-28)}{2*7}\)

Mehek (mehek14):

\(\dfrac{28}{14}=2\)

OpenStudy (anonymous):

28/14=2

Mehek (mehek14):

the plug 2 into the equation for x

OpenStudy (anonymous):

oh u did it before i could lol

OpenStudy (anonymous):

2 is x?

Mehek (mehek14):

\(f(2)=7*2^2 - 28*2 - 60\)

Mehek (mehek14):

yes x=2

OpenStudy (anonymous):

-88?

Mehek (mehek14):

correct

OpenStudy (anonymous):

2,-88

Mehek (mehek14):

yup

OpenStudy (anonymous):

good to see u back on too lol

Mehek (mehek14):

:)

OpenStudy (anonymous):

oh no, i dnt get the next one at all...post here or new post?

Mehek (mehek14):

new?

OpenStudy (anonymous):

wait i might be able to do it, let me post it and show my work before u comment to check it?

Mehek (mehek14):

ok